Check, whether the collection G, given by

G' =. { ] 1/(n+2), 1/n [ : n ∈ N}

is an open cover of ]0,1[


1
Expert's answer
2021-06-22T09:02:26-0400

(0,1) has no finite subcover because,if possible it has a subcover

i.e then "U^n_{i=1}I_n \u2265(0,1)"

"I_n=(\frac{1}{n+2}, \frac{1}{n})"

But "\frac{1}{m+1}< \frac{1}{n}" and "0<\frac{1}{m+1}<1"

But "\frac{1}{m+1} \notin U^n_{i=1}I_n" So, "[I_n ]_{n=1}^m" is not a subcover

That is to say for every m , we see that "\frac{1}{m+1} \notin U_{n=1}^m I_n"

So it has no subcover